# Non-squeezing and other global rigidity results in locally conformal symplectic geometry

### v1

@@ 2. Locally conformal symplectic manifolds
In this section we present the basic notions and results on lcs geometry that are needed in the rest of the article. Although the material is classic (except for 2.11 on the lcs product and graph, introduced by Chantraine and Sackel [CS24], and 2.10 on orderability), we include detailed proofs whenever we could not find them in the literature. Among the basic results discussed in this section, Corollary 2.9 is particularly important for us, because it allows to define the (unique) lift to the symplectic cover of a divergence free lcs diffeomorphism. Using this notion we give in Section 2.11 a description of the lcs graph of a divergence free lcs diffeomorphism that seems simpler than the one in [CS24]. In 2.10 we introduce the notion of orderability in lcs geometry, and prove Theorem 1.3 (i).
@@ 2.5. Lcs vector fields and isotopies
for a $1$-parameter family of real numbers $a_{t}$. Denote by $\bar{X}_{t}$ the vector field generating $\{\bar{\varphi}_{t}\}$. Using Lemmas 2.14 and 2.15 we obtain
@@ 2.6. Lcs Hamiltonian isotopies
By Lemma 2.16, $X_{t}$ is divergence free and so $d_{\eta}(\iota_{X_{t}}\omega)=0$ for all $t$. By Lemmas 2.14 and 2.15 we thus have
@@ 2.9. Contactization
Moreover, by Lemmas 2.14 and 2.15 we have $\frac{dg_{t}}{dt}=\eta(X_{t})\circ\varphi_{t}$. We thus obtain
@@ 2.10. Orderability
Similarly, for a connected lcs manifold $\big(M,[(\eta,\omega)]\big)$ such that $\eta$ is not exact we consider the partial relation $\leq$ on the universal cover of the group of compactly supported lcs Hamiltonian diffeomorphisms defined by posing $\{\varphi_{t}^{(1)}\}]\leq[\{\varphi_{t}^{(2)}\}]$ if $\{\varphi_{t}^{(2)}\}]\cdot[\{\varphi_{t}^{(1)}\}]^{-1}$ can be represented by a non-negative lcs Hamiltonian isotopy. We say that $\big(M,[(\eta,\omega)]\big)$ is orderable if $\leq$ is a bi-invariant partial order. As in the contact case, reflexivity, transitivity and bi-invariance of $\leq$ are always true, and can be easily proved using Lemmas 2.20 and 2.21. On the other hand, anti-symmetry is equivalent to the non-existence of a non-negative non-constant contractible loop of compactly supported lcs Hamiltonian diffeomorphisms, and can fail in general. For instance, by Example 2.19 any non-negative non-constant contractile loop of contactomorphisms on a contact manifold lifts to a non-negative non-constant contractile loop of lcs Hamiltonian diffeomorphisms of the locally conformal symplectization. Thus, the locally conformal symplectization of a contact manifold that is not orderable (for instance the standard contact sphere [EKP06]) is also not orderable. The converse implication is in general not clear, but we will see in Section 7 that it holds for the locally conformal symplectizations of $(\mathbb{R}^{2n+1},\xi_{0})$ and $(\mathbb{R}^{2n}\times S^{1},\xi_{0})$.
@@ 3. Essential Lee chords and translated points
thus if and only if $(\varphi_{f(q)}^{\omega})_{\ast}\big(T_{d_{\beta}f(q)}L_{f,\beta}\big)\cap T_{0_{q}}0_{B}\neq\{0\}$, and so if and only if there is $Y\in T_{q}B$ such that $\big(\varphi_{f(q)}^{\omega}\circ d_{\beta}f\big)_{\ast}\,(Y)\in T_{0_{q}}0_{B}$. Consider local coordinates $x=(x_{1},\cdots,x_{n})$ on a neighborhood of $q$, and write $Y=\sum_{j=1}^{n}Y_{x_{j}}\frac{\partial}{\partial x_{j}}$ and $Y_{x}=\left(\begin{smallmatrix}Y_{x_{1}}\\ \cdots\\ Y_{x_{n}}\end{smallmatrix}\right)$. Also write $\beta=\sum_{j=1}^{n}\beta_{j}\,dx_{j}$. In the associated coordinates on $T^{\ast}B$ we then have
$d_{\beta}f(x)=\Big(x\,,\,\frac{\partial f}{\partial x_{1}}(x)-f(x)\,\beta_{1}\,,\,\cdots\,,\,\frac{\partial f}{\partial x_{n}}(x)-f(x)\,\beta_{n}\Big)\,.$
Since $q$ is a critical point of $f$, we have
$(d_{\beta}f)_{\ast}(Y)=\big(Y_{x}\,,\,\operatorname{Hess}_{q}(f)\cdot Y_{x}-f(q)\,d\beta\cdot Y_{x}\big)\,$,
where we write $d\beta=\big(\frac{\partial\beta_{i}}{\partial x_{j}}\big)_{ij}$. Thus
$(\varphi_{f(q)}^{\omega}\circ d_{\beta}f)_{\ast}\,(Y)=\big(Y_{x}\,,\,\operatorname{Hess}_{q}(f)\cdot Y_{x}\big)\,$,
and so $(\varphi_{f(q)}^{\omega}\circ d_{\beta}f)_{\ast}\,(Y)\in T_{0_{q}}0_{B}$ if and only if $\operatorname{Hess}_{q}(f)\cdot Y_{x}=0$. This shows that $\big(f(q),\gamma_{q}\big)$ is not transverse if and only if the Hessian of $f$ at $q$ is degenerate, thus if and only if $q$ is a degenerate critical point of $f$. ∎
@@ 4. Generating functions for Lagrangian submanifolds of twisted cotangent bundles
thus if and only if $(\varphi_{F(e)}^{\omega})_{\ast}\big(T_{i_{F,\beta}(e)}L\big)\cap T_{0_{q}}0_{B}\neq\{0\}$, and so if and only if there is $Y\in T_{e}\Sigma_{F}$ such that $(\varphi_{F(e)}^{\omega}\circ i_{F,\beta})_{\ast}\,(Y)\in T_{0_{q}}0_{B}$. Consider local coordinates $x=(x_{1},\cdots,x_{n})$ on a neighborhood $\mathcal{U}$ of $q$ in $B$, and coordinates $(x,\zeta)=(x_{1},\cdots,x_{n},\zeta_{1},\cdots,\zeta_{N})$ on $\mathcal{U}\times\mathbb{R}^{N}$. Write
$Y=\sum_{j=1}^{n}Y_{x_{j}}\frac{\partial}{\partial x_{j}}+\sum_{j=1}^{N}Y_{\zeta_{j}}\frac{\partial}{\partial\zeta_{j}}\,$,
$Y_{x}=\left(\begin{smallmatrix}Y_{x_{1}}\\ \cdots\\ Y_{x_{n}}\end{smallmatrix}\right)$ and $Y_{\zeta}=\left(\begin{smallmatrix}Y_{\zeta_{1}}\\ \cdots\\ Y_{\zeta_{N}}\end{smallmatrix}\right)$. In the associated coordinates on $T^{\ast}B$ we then have
$(i_{F,\beta})_{\ast}(Y)=\Big(Y_{x}\,,\,\frac{\partial^{2}F}{\partial x^{2}}(e)\cdot Y_{x}+\frac{\partial^{2}F}{\partial x\partial\zeta}(e)\cdot Y_{\zeta}-F(e)\,d\beta\cdot Y_{x}\Big)\,$,
where we write $\beta=\sum_{j=1}^{n}\beta_{j}dx_{j}$ and $d\beta=\big(\frac{\partial\beta_{i}}{\partial x_{j}}\big)_{ij}$. We thus have
$(\varphi_{F(e)}^{\omega}\circ i_{F,\beta})_{\ast}(Y)=\Big(Y_{x}\,,\,\frac{\partial^{2}F}{\partial x^{2}}(e)\cdot Y_{x}+\frac{\partial^{2}F}{\partial x\partial\zeta}(e)\cdot Y_{\zeta}\Big)$
and so $(\varphi_{F(e)}^{\omega}\circ i_{F,\beta})_{\ast}(Y)\in T_{0_{q}}0_{B}$ if and only if
$\frac{\partial^{2}F}{\partial x^{2}}(e)\cdot Y_{x}+\frac{\partial^{2}F}{\partial x\partial\zeta}(e)\cdot Y_{\zeta}=0\,.$
Since $Y\in T_{e}\Sigma_{F}$ we also have
$\frac{\partial^{2}F}{\partial x\partial\zeta}(e)\cdot Y_{x}+\frac{\partial^{2}F}{\partial\zeta^{2}}(e)\cdot Y_{\zeta}=0\,.$
We conclude that $\big(F(e),\gamma_{e}\big)$ is not transverse if and only if there is $Y\in T_{e}E$ such that $\operatorname{Hess}_{e}(F)\cdot Y=0$, and so if and only if $e$ is a degenerate critical point of $F$. ∎
@@
    Non-degeneracy: if $F:B\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a generating function quadratic at infinity for a Lagrangian submanifold $L$ of $T^{\ast}B$ and if $c([B],F)=c(1,F)$ then $L$ is the zero section.
@@
We have to show that the relation $\leq$ on $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n+1})$ or $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ is a bi-invariant partial order. Reflexivity follows from the fact that $\{\operatorname{id}\}$ has Hamiltonian function $H_{t}\equiv 0$. Transitivity and bi-invariance follow from Lemmas 2.20 and 2.21. As for anti-symmetry, if $\varphi_{1}\leq\varphi_{2}$ and $\varphi_{2}\leq\varphi_{1}$ then by Proposition 7.2 we also have $\varphi_{1}\preceq\varphi_{2}$ and $\varphi_{2}\preceq\varphi_{1}$. But, by Theorem 7.1, $\preceq$ is anti-symmetric. We thus conclude that $\varphi_{1}=\varphi_{2}$. ∎
@@
We proceed exactly as in [SS24]. Suppose by contradiction that there exists $g\in\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ whose conjugacy class is dense in $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$. Since $g$ has compact support, we can assume that its support is contained in $S^{1}\times B(R)\times S^{1}$ for some $R$. Choose $f\in\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ such that $f(x)\neq x$ for all $x\in S^{1}\times B(R+1)\times S^{1}$. Such $f$ can be found by considering the time-$1$ map of a lcs Hamiltonian isotopy with Hamiltonian function that is equal to $\frac{1}{2}$ on $S^{1}\times B(R+1)\times S^{1}$ and vanishes outside a larger set. Since we assume that the conjugacy class of $g$ is dense in $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$, there exists a sequence $(h_{n})$ in $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ such that $(h_{n}\circ g\circ h_{n}^{-1})$ converges to $f$. This implies that
@@
for all $n$. Indeed, if $S^{1}\times B(R+1)\times S^{1}\subset h_{n}(\operatorname{supp}(g))$ then $S^{1}\times B(R+1)\times S^{1}\subset h_{n}(S^{1}\times B(R)\times S^{1})$, and so $h_{n}^{-1}(S^{1}\times B(R+1)\times S^{1})\subset S^{1}\times B(R)\times S^{1}$, contradicting Theorem 1.4. By taking $n$ large enough in (19) we obtain
$S^{1}\times B(R+1)\times S^{1}\not\subset\operatorname{supp}(f)\,$,
and thus there exists $x\in S^{1}\times B(R+1)\times S^{1}$ such that $f(x)=x$, contradicting the construction of $f$. ∎

### v2

@@ 2. Locally conformal symplectic manifolds
In this section we present the basic notions and results on lcs geometry that are needed in the rest of the article, and take the opportunity to also discuss $\mathcal{C}^{0}$-rigidity of the group of lcs diffeomorphisms 2.5. Although the material is classic (except for $\mathcal{C}^{0}$-rigidity 2.5, for the lcs product and graph 2.12, introduced by Chantraine and Sackel [CS24], and for orderability 2.11), we include detailed proofs whenever we could not find them in the literature. Among the basic results discussed in this section, Corollary 2.9 is particularly important for us, because it allows to define the (unique) lift to the symplectic cover of a divergence free lcs diffeomorphism. Using this notion we give in Section 2.12 a description of the lcs graph of a divergence free lcs diffeomorphism that seems simpler than the one in [CS24]. In 2.11 we introduce the notion of orderability in lcs geometry, and prove Theorem 1.3 (i).
@@ 2.4. Lcs diffeomorphisms
2.5. $\mathcal{C}^{0}$-rigidity of lcs diffeomorphisms
The group of lcs diffeomorphisms of a lcs manifold $(M,[(\eta,\omega)])$ is $\mathcal{C}^{0}$-closed in the group of diffeomorphisms of $M$. As explained hereafter, this is a direct consequence of the Eliashberg–Gromov $\mathcal{C}^{0}$-rigidity of symplectomorphisms [Eli87]. Let us fix an auxiliary Riemannian metric $g$ on $M$.
Proposition 2.14.
Let $(\varphi_{k})_{k\in\mathbb{N}}$ be a sequence of lcs diffeomorphisms of a lcs manifold $(M,[(\eta,\omega)])$ that converges in the $\mathcal{C}^{0}$-topology to a diffeomorphism $\varphi$ of $M$. Then $\varphi$ is a lcs diffeomorphism of $(M,[(\eta,\omega)])$.
Proof.
Let $p$ be a point of $M$, and suppose first that $\varphi(p)=p$ and $\varphi_{k}(p)=p$ for all $k\in\mathbb{N}$. Let $\mathcal{U}$ be a contractible open neighborhood of $p$ in $M$ whose closure is compact and whose boundary is a smooth submanifold of $M$. Set $\mathcal{V}_{k}=\varphi_{k}(\mathcal{U})$ and $\mathcal{V}=\varphi(\mathcal{U})$. Let $\mathcal{W}$ denote a contractible open neighborhood of $p$ that contains $\mathcal{V}\cup_{k}\mathcal{V}_{k}$ and whose closure is compact as well. Let $f\in\mathcal{C}^{\infty}(\mathcal{W})$ be such that $\eta|_{\mathcal{W}}=df$ and set $\overline{\omega}=e^{-f}\omega$. We may assume without loss of generality that $\mathcal{W}$ is the domain of a Darboux chart $(\mathcal{W},\Psi)$ with $\Psi(\mathcal{W})=B^{2n}(R)$, $\Psi(\mathcal{U})=B^{2n}(r)$ (and $\Psi^{*}\omega_{0}=\overline{\omega}$). Now, we have $\varphi_{k}^{*}(\overline{\omega})=e^{c_{k}}\overline{\omega}$ for some $c_{k}\in\mathbb{R}$. Notice that the $\mathcal{C}^{0}$-convergence of $(\varphi_{k})_{k\in\mathbb{N}}$ towards $\varphi$ implies that the symplectic volume of $\varphi_{k}(\mathcal{U})$ converges towards the symplectic volume of $\varphi(\mathcal{U})$, or in other words that
(1) $\lim_{k\to\infty}\int_{\varphi_{k}(\mathcal{U})}\overline{\omega}^{n}=\int_{\varphi(\mathcal{U})}\overline{\omega}^{n}.$
Indeed, for $\varepsilon>0$, let $N_{\varepsilon}\subset\mathcal{W}$ denote the image of a tubular neighborhood of $\partial\varphi(\mathcal{U})$ such that
$\int_{N_{\varepsilon}}\overline{\omega}^{n}<\varepsilon\,.$
This exists because $\partial\varphi(\mathcal{U})$ is a closed embedded submanifold of $\mathcal{W}$, whose Lebesgue measure is 0. Besides, there exists $K\in\mathbb{N}$ such that for all $k\geq K$, $\varphi_{k}(\partial\mathcal{U})\subset N_{\varepsilon}$, which implies that
$\left|\int_{\varphi_{k}(\mathcal{U})}\overline{\omega}^{n}-\int_{\varphi(\mathcal{U})}\overline{\omega}^{n}\right|<\varepsilon\,.$
The relation (1) is equivalent to
$\int_{\mathcal{U}}\varphi^{*}(\overline{\omega}^{n})=\lim_{k\to\infty}\int_{\mathcal{U}}\varphi_{k}^{*}(\overline{\omega}^{n})=\lim_{k\to\infty}e^{nc_{k}}\int_{\mathcal{U}}\overline{\omega}^{n}\,.$
Hence the sequence $(c_{k})_{k\in\mathbb{N}}$ has a limit $c$ in $\mathbb{R}$ (if the group of periods of $\eta$ had rank one, or, more generally, was discrete, we could argue that there exists a $K\in\mathbb{N}$ such that $c_{k}=c_{K}$ for all $k\geq K$, but this is not necessarily the case). Set
$\psi_{k}:\mathbb{R}^{2n}\to\mathbb{R}^{2n}:x\mapsto e^{-\frac{1}{2}c_{k}}x$
and
$\psi:\mathcal{W}\to\mathcal{W}:x\mapsto e^{-\frac{1}{2}c}x\,.$
This says that $\psi_{k}\circ\varphi_{k}:(\mathcal{U},\overline{\omega}|_{\mathcal{U}})\to(\mathcal{W},\overline{\omega})$ is a sequence of symplectic embeddings that $C^{0}$-converges to the embedding $\psi\circ\varphi:(\mathcal{U},\overline{\omega})\to(\mathcal{W},\overline{\omega})$. The $C^{0}$-rigidity of symplectic embeddings implies that $\psi\circ\varphi$ is a symplectic embedding, hence that $\varphi$ is a conformally symplectic embedding.
[…]
@@ 2.6. Lcs vector fields and isotopies
for a $1$-parameter family of real numbers $a_{t}$. Denote by $\bar{X}_{t}$ the vector field generating $\{\bar{\varphi}_{t}\}$. Using Lemmas 2.15 and 2.16 we obtain
@@ 2.7. Lcs Hamiltonian isotopies
By Lemma 2.17, $X_{t}$ is divergence free and so $d_{\eta}(\iota_{X_{t}}\omega)=0$ for all $t$. By Lemmas 2.15 and 2.16 we thus have
@@ 2.10. Contactization
Moreover, by Lemmas 2.15 and 2.16 we have $\frac{dg_{t}}{dt}=\eta(X_{t})\circ\varphi_{t}$. We thus obtain
@@ 2.11. Orderability
Similarly, for a connected lcs manifold $\big(M,[(\eta,\omega)]\big)$ such that $\eta$ is not exact we consider the partial relation $\leq$ on the universal cover of the group of compactly supported lcs Hamiltonian diffeomorphisms defined by posing $\{\varphi_{t}^{(1)}\}]\leq[\{\varphi_{t}^{(2)}\}]$ if $\{\varphi_{t}^{(2)}\}]\cdot[\{\varphi_{t}^{(1)}\}]^{-1}$ can be represented by a non-negative lcs Hamiltonian isotopy. We say that $\big(M,[(\eta,\omega)]\big)$ is orderable if $\leq$ is a bi-invariant partial order. As in the contact case, reflexivity, transitivity and bi-invariance of $\leq$ are always true, and can be easily proved using Lemmas 2.21 and 2.22. On the other hand, anti-symmetry is equivalent to the non-existence of a non-negative non-constant contractible loop of compactly supported lcs Hamiltonian diffeomorphisms, and can fail in general. For instance, by Example 2.20 any non-negative non-constant contractile loop of contactomorphisms on a contact manifold lifts to a non-negative non-constant contractile loop of lcs Hamiltonian diffeomorphisms of the locally conformal symplectization. Thus, the locally conformal symplectization of a contact manifold that is not orderable (for instance the standard contact sphere [EKP06]) is also not orderable. The converse implication is in general not clear, but we will see in Section 7 that it holds for the locally conformal symplectizations of $(\mathbb{R}^{2n+1},\xi_{0})$ and $(\mathbb{R}^{2n}\times S^{1},\xi_{0})$.
@@ 3. Essential Lee chords and translated points
thus if and only if $(\varphi_{f(q)}^{\omega})_{\ast}\big(T_{d_{\beta}f(q)}L_{f,\beta}\big)\cap T_{0_{q}}0_{B}\neq\{0\}$, and so if and only if there is $Y\in T_{q}B$ such that $\big(\varphi_{f(q)}^{\omega}\circ d_{\beta}f\big)_{\ast}\,(Y)\in T_{0_{q}}0_{B}$. Write $Y=\dot{\delta}(t)$ for a path $\delta:(-\varepsilon,\varepsilon)\to B$ with $\delta(0)=q$. Then
$\big(\varphi^{\omega}_{f(q)}\circ d_{\beta}f\big)_{*}(Y)=$ $\left.\frac{d}{dt}\right\lvert_{t=0}\;\varphi^{\omega}_{f(q)}\circ d_{\beta}f\,\big(\delta(t)\big)$
$=$ $\left.\frac{d}{dt}\right\lvert_{t=0}\;df\big(\delta(t)\big)+\Big(f(q)-f\big(\delta(t)\big)\Big)\,\beta\big(\delta(t)\big)$
$=$ $\;Y+d^{2}f\,(Y,\;\cdot\;)\,$,
where the symbol ++ in the last equality refers to the splitting of $T_{0_{q}}T^{*}B$ into horizontal and vertical subspaces, and where we have used that $q$ is a critical point of $f$. This shows that $\big(f(q),\gamma_{q}\big)$ is not transverse if and only if the Hessian of $f$ at $q$ is degenerate, thus if and only if $q$ is a degenerate critical point of $f$. ∎
@@ 4. Generating functions for Lagrangian submanifolds of twisted cotangent bundles
thus if and only if $(\varphi_{F(e)}^{\omega})_{\ast}\big(T_{i_{F,\beta}(e)}L\big)\cap T_{0_{q}}0_{B}\neq\{0\}$, and so if and only if there is $Y\in T_{e}\Sigma_{F}$ such that
(8) $(\varphi_{F(e)}^{\omega}\circ i_{F,\beta})_{\ast}\,(Y)\in T_{0_{q}}0_{B}\,.$
Write $Y=\dot{\delta}(t)$ for a path $\delta:(-\varepsilon,\varepsilon)\to\Sigma_{F}$ with $\delta(0)=e$. Set $\delta_{o}=\operatorname{pr}_{1}\circ\,\delta$, where $\operatorname{pr}_{1}:\Sigma_{F}\subset E=B\times\mathbb{R}^{N}\to B$ is the projection to the first factor. Observe that $N_{E}^{*}$ is canonically isomorphic to the pullback bundle $\operatorname{pr}_{1}^{\,\ast}\,T^{*}B$, and denote by $p:N_{E}^{*}\to T^{*}B$ the natural projection. We can then write $i_{F,\beta}(e)=p\,(d_{\beta}F(e))$ and
$(\varphi_{F(e)}^{\omega}\circ i_{F,\beta})_{\ast}\,(Y)$ $=\left.\frac{d}{dt}\right\lvert_{t=0}p\,\Big(d_{\beta}F\big(\delta(t)\big)\Big)+F(e)\,\beta\big(\delta_{o}(t)\big)$
$=\left.\frac{d}{dt}\right\lvert_{t=0}p\,\Big(dF\big(\delta(t)\big)\Big)+\Big(F(e)-F\big(\delta(t)\big)\Big)\,\beta\big(\delta_{o}(t)\big)\,.$
We observe that $\delta_{1}(t):=p\,\left(dF(\delta(t))\right)$ and $\delta_{2}(t):=\big(F(e)-F(\delta(t))\big)\,\beta(\delta_{o}(t))$ are two paths in $T^{*}B$ with $\delta_{1}(0)=\delta_{2}(0)=0_{q}$. Therefore, their tangent vectors at $t=0$ both belong to $T_{0_{q}}T^{*}B$. The latter vector space admits a canonical decomposition as the direct sum of $T_{q}B$ – the “horizontal” part – and $T^{*}_{q}B$ – the “vertical” part. The decompositions of $\left.\frac{d}{dt}\right\lvert_{t=0}\delta_{1}(t)$ and $\left.\frac{d}{dt}\right\lvert_{t=0}\delta_{2}(t)$ into horizontal and vertical parts are
$\left.\frac{d}{dt}\right\lvert_{t=0}\delta_{1}(t)=p_{*}\big(Y+d^{2}F(Y,\,\cdot\,)\big)={\operatorname{pr}_{1}}_{*}(Y)+p_{\ast}\big(d^{2}F(Y,\,\cdot\,)\big)$
and
$\left.\frac{d}{dt}\right\lvert_{t=0}\delta_{2}(t)={\operatorname{pr}_{1}}_{*}(Y)+\left.\frac{d}{dt}\right\lvert_{t=0}\Big(F(e)-F\big(\delta(q)\big)\Big)\,\beta(q)={\operatorname{pr}_{1}}_{*}(Y)+0_{q}\,$,
where we have used that $e$ is a critical point of $F$. The condition (8) is equivalent to asking that the sum of the vertical parts of $\left.\frac{d}{dt}\right\lvert_{t=0}\delta_{1}(t)$ and $\left.\frac{d}{dt}\right\lvert_{t=0}\delta_{2}(t)$ vanishes. Since this sum is $p_{\ast}\big(d^{2}F(Y,\,\cdot\,)\big)$, we conclude that the Lee chord $\big(F(e),\gamma_{e}\big)$ is not transverse if and only if there is $Y\in T_{e}\Sigma_{F}$ such that $p_{\ast}\big(d^{2}F(Y,\,\cdot\,)\big)$ vanishes, hence $d^{2}F(Y,\,\cdot\,)$ vanishes, and so if and only if $e$ is a degenerate critical point of $F$. ∎
@@
    Non-degeneracy: if $F:B\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a generating function quadratic at infinity such that $c([B],F)=c(1,F)$ then $F$ generates the zero section of $T^{\ast}B$.
@@
We have to show that the relation $\leq$ on $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n+1})$ or $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ is a bi-invariant partial order. Reflexivity follows from the fact that $\{\operatorname{id}\}$ has Hamiltonian function $H_{t}\equiv 0$. Transitivity and bi-invariance follow from Lemmas 2.21 and 2.22. As for anti-symmetry, if $\varphi_{1}\leq\varphi_{2}$ and $\varphi_{2}\leq\varphi_{1}$ then by Proposition 7.2 we also have $\varphi_{1}\preceq\varphi_{2}$ and $\varphi_{2}\preceq\varphi_{1}$. But, by Theorem 7.1, $\preceq$ is anti-symmetric. We thus conclude that $\varphi_{1}=\varphi_{2}$. ∎
@@
We proceed exactly as in [SS24]. Suppose by contradiction that there exists $g\in\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ whose conjugacy class is dense in $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$. Since $g$ has compact support, we can assume that its support is contained in $S^{1}\times B(R)\times S^{1}$ for some $R$. Choose $f\in\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ such that $f(x)\neq x$ for all $x\in S^{1}\times B(R+1)\times S^{1}$. Such $f$ can be found by considering the time-$1$ map of a lcs Hamiltonian isotopy with Hamiltonian function that is equal to $\frac{1}{2}$ on $S^{1}\times B(R+1)\times S^{1}$ and vanishes outside a larger set. Since we assume that the conjugacy class of $g$ is dense in $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$, there exists a sequence $(h_{n})$ in $\operatorname{Ham}^{c}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ such that $(h_{n}\circ g\circ h_{n}^{-1})$ converges to $f$. Theorem 1.4 implies that
@@
for all $n$. Indeed, if, on the contrary, $S^{1}\times B(R+1)\times S^{1}\subset h_{n}(\operatorname{supp}(g))$ then $S^{1}\times B(R+1)\times S^{1}\subset h_{n}(S^{1}\times B(R)\times S^{1})$, and so $h_{n}^{-1}(S^{1}\times B(R+1)\times S^{1})\subset S^{1}\times B(R)\times S^{1}$, contradicting Theorem 1.4.
Now 21 implies the existence of a sequence $(z_{n})_{n\in\mathbb{N}}$ in $S^{1}\times B(R+1)\times S^{1}$ such that
$z_{n}=h_{n}\circ g\circ h_{n}^{-1}(z_{n}).$
Since $S^{1}\times B(R+1)\times S^{1}$ is compact, the sequence $(z_{n})$ admits a converging subsequence $(z_{n_{k}})_{k\in\mathbb{N}}$. Its limit $z\in S^{1}\times B(R+1)\times S^{1}$ is necessarily a fixed point of $f$, contradicting the construction of the latter function.